# A Neural-Guided Variational Quantum Algorithm for Efficient Sign Structure Learning in Hybrid Architectures

### v1

@@ 3. Scaling Under Sampling Errors
Figure 5: Scaling test of sVQNHE using the MaxCut problem as an example. (Left) The relative solution quality for the MaxCut problem, comparing parameter sets 1 and 2 across different problem sizes. The methods sign-VQNHE, brickwork-VQE, and sign-VQE correspond to the sVQNHE method with a 2-layer sign ansatz, the VQE method with brickwork ansatz, and the VQE method with 2-layer sign ansatz, respectively. The depth of brickwork ansatz is determined according to the reference [36], that is, the circuit depth is sublinearly related to the number of vertices in the graph. For sign-VQE, the ansatz is the same as the circuit finally generated by the corresponding sVQNHE scheme finally generated, utilizing a two-layer sign ansatz with two diagonal and two non-diagonal layers. The value range of relative solution quality is $[0,1]$. The closer to 1, the better the performance is compared with other algorithms (see Equation (14) in the Method section for details). For example, in the case of 45 vertices, the quality of the solution of sVQNHE is relatively optimal regardless of whether the parameter set is 1 or 2. (Right) The single-step measurement consumption for the same problem sizes. Parameter set 2 features an increased number of parameters in the NN component and doubles the number of iterations compared to parameter set 1. The circuit depth in the brickwork architecture scales sub-linearly with the number of vertices.
@@ Discussion
We present sVQNHE, a hybrid quantum-classical framework that strategically optimizes the roles of quantum and classical resources. Compared to the original VQNHE, our approach significantly improves measurement efficiency, substantially suppresses the numerical instability of the neural network, and enhances overall convergence.
Notably, sVQNHE leverages commuting diagonal gates to encode quantum sign structures, allowing efficient gradient estimation with far fewer measurements. The classical neural network is tasked with learning wavefunction amplitudes and guiding parameter updates, effectively sidestepping the barren plateau issue commonly seen in deep quantum circuits. This division of labor enables a layer-wise optimization strategy that remains stable even under noise and sampling limitations.
Our numerical simulations demonstrate compelling improvements across a broad class of computational problems. In the 6-qubit J1-J2 model, the sVQNHE method with a single-layer assumption achieves more than 7-fold the average accuracy and 19-fold the convergence speed of the 2-layer hardware-efficient VQE. Compared with the baseline NN, the average accuracy is improved by nearly two orders of magnitude. In sign-free models such as the TFIM and Ising chains, sVQNHE consistently outperforms QAOA and VQE, achieving broader expressivity and better scalability. For combinatorial optimization, such as MaxCut on 45-node graphs, sVQNHE improves solution quality by $19\%$ and quantum resource efficiency by $85\%$ compared to brickwork ansatz-based VQE.
Taken together, these results highlight sVQNHE’s broad applicability, efficiency, and scalability. By tightly integrating neural networks with quantum circuits along with a revised role that each component plays in the new algorithm, sVQNHE offers a practical path toward quantum utility on NISQ and early fault-tolerant devices. Looking ahead, this framework can be further extended by incorporating advanced machine learning models and better measurement strategies, and applied to a wider class of many-body simulation and large-scale optimization problems. Extending sVQNHE to tackle PQC-based quantum machine learning is also conceptually straightforward. As quantum hardware continues to evolve, we expect sVQNHE to remain a strong candidate for bridging theoretical capabilities with real-world impact.
@@ Methods
(Expressiveness of the Quantum Circuit Ansatz)[40, 41, 12, 42, 43] Let $\mathfrak{g_{1}}$ and $\mathfrak{g_{2}}$ be the Lie algebras corresponding to the generators $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, where:
@@
$\gradient_{\theta}f(x;\theta)=\frac{1}{2}[f(x;\theta+\frac{\pi}{2})-f(x;\theta-\frac{\pi}{2})].$ (8)
@@
Since $W_{k}$ and $F$ are all commutative, and $E_{P_{i}}(\theta^{W})=\bra{\psi_{0}}W^{\dagger}(\theta^{W})F^{\dagger}P_{i}FW(\theta^{W})\ket{\psi_{0}}$, we have
@@
$=\bra{\psi_{0}}W^{\dagger}(\theta^{W})F^{\dagger}W_{k}^{\dagger}(\pm\frac{\pi}{2})P_{i}W_{k}(\pm\frac{\pi}{2})FW(\theta^{W})\ket{\psi_{0}}.$ (10)
@@
In summary, due to the commutation properties of the Sign Ansatz, our algorithm requires fewer measurements compared to general algorithms that rely on parameterized quantum circuits. For instance, considering the 1D J1-J2 model presented below (11), if $|p|\geq 2$ in the Sign Ansatz, the total number of measurements in a single iteration of our algorithm is equal to $13(n_{1}+n_{2})+$ 1, where $n_{1}$ is the number of $\langle i,j\rangle$, and $n_{2}$ is the number of $\langle\langle i,j\rangle\rangle$.
@@
We use TensorCircuit[48] to efficiently obtain all the quantum circuits and the neural network numerical results with the following Ansatz setting, and the details of the VQNHE algorithm are in the TensorCircuit Demo.
[…]
The Sign Ansatz design equivalents for the sampling and noise versions of the numerical results with finite shots are detailed as follows. A combination of $YM$, $YP$, $C_{z}$, and $R_{z}$ gates equivalently replaces the $R_{zz}$ gate. Similarly, the $R_{y}$ gate is equivalently replaced by a combination of $XP$, $XM$, and $R_{z}$ gates. Here, $YM$ and $YP$ denote positive and negative rotations by $\frac{\pi}{4}$ around the Y-axis of the Bloch sphere, respectively, while $XM$ and $XP$ denote positive and negative rotations by $\frac{\pi}{4}$ around the X-axis of the Bloch sphere, respectively. The two-qubit gates used in our cases connect the edges of the model (1D J1-J2 model and 2D Heisenberg model).
[…]
In the design of Hardware Efficient Ansatz, where the $CNOT$ is only connecting the two nearest qubits (2D Heisenberg Model), n is the number of repetitions. Note that after repeating the above block n times, one more layer of $R_{y}$ and one more layer of $R_{z}$ are added. For example, for HEA2 used in this paper, there are three $R_{y}$ layers, three $R_{z}$ layers, and two $CNOT$ layers. Similarly, the $CNOT$ gate in the numerical simulation of the finite-shot version is equivalently replaced by a combination of $YM$, $YP$, and $C_{z}$ gates.
Metrics
In the experiment, we use the following metrics to evaluate the performance of the method:
  • Mean absolute error ratio $R_{MAE}$:
    $R_{MAE}=\frac{MAE_{1}-MAE_{0}}{MAE_{0}}.$ (12)
    The metric $R_{MAE}$ represents the relative reduction in mean absolute error (MAE) of different algorithms compared to the baseline algorithm (NN) with the same number of iterations. Here $MAE_{0}$ represents the MAE of the baseline algorithm (NN) being compared. The smaller the $R_{MAE}$ value, the better the compared algorithm is relative to NN.
  • Variance ratio $R_{Var}$:
    $R_{Var}=\frac{Var_{1}-Var_{0}}{Var_{0}}.$ (13)
    The metric $R_{Var}$ represents the relative reduction in the variance of different algorithms compared to the baseline algorithm (NN) given the same number of iterations. Among them, $Var_{0}$ represents the variance of the baseline algorithm (NN) being compared. The smaller the $R_{Var}$ value, the more stable the compared algorithm is relative to NN, and the smaller the variability of the results.
  • Estimated approximation ratio $R_{e}$ for the MaxCut problem:
    $R_{e}=\frac{MaxCut_{algo}}{MaxCut_{max}}.$ (14)
    The metric $R_{e}$ provides an estimated approximation ratio based on the best-known result for the MaxCut problem, which measures the relative solution quality. Here, $MaxCut_{max}$ represents the maximum MaxCut value achieved across all evaluated algorithms, while $MaxCut_{algo}$ denotes the MaxCut value obtained by the specific algorithm under assessment. A value of $R_{e}$ approaching $100\%$ indicates superior performance relative to the other algorithms compared.
Acknowledgements
CYH acknowledges support from National Ke Research and Development Program of China (2024FA1306400), Natural Science Foundation of China (22373085). AH gratefully acknowledges the sponsorship from Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. CityU 11200120), City University of Hong Kong (Project No. 7005615, 7006103), CityU Seed Fund in Microelectronics (Project No. 9229135), and Hon Hai Research Institute (Project No. 9231594). YYS acknowledges support from the National Natural Science Foundation of China (Grant No. 12404575), the Shandong Provincial Natural Science Foundation (Grant No. ZR2022LLZ008), the Natural Science Foundation of Shanghai (Grant No. 23ZR1469600), and the Shanghai Sailing Program (Grant No. 23YF1452600).
Code and data availability
The source code and data associated with this project are publicly available. They can be accessed at the following GitHub repository: https://github.com/renmengzhen/sVQNHE
Supplementary
Optimizing Quantum Resources with Sign Ansatz VQNHE
In this section, we evaluate the quantum resource efficiency of the sVQNHE framework against the Hardware Efficient Ansatz VQE. First, we test different ansatz on the 9-qubit 2D Heisenberg model with the initial state $\ket{++...++}$. The results are shown in Figure 6, it can be seen that sVQNHE uses very few quantum resources to converge to the desired accuracy compared to both neural networks and VQE, which implies that the sVQNHE reaches the desired accuracy with fewer quantum resources compared to VQE in the number of circuits, and the number of parameters. Besides, since the zero quantum resource results indicate that without the assistance of Sign Ansatz the neural network itself cannot reach good accuracy, we can secondly claim that this framework can greatly improve the expressiveness of the neural network with limited quantum resources. In addition, we also verified the quantum resource efficiency of sVQNHE relative to VQE on the J1-J2 model. The results are shown in the Table 3.
Figure 6: 9-qubit Heisenberg Model Result. The x-axis shows sorted results by the converged energy with different random initial parameters. The y-axis is the energy difference from the ground state. The red, green, and blue dashed lines are responding to Hardware Efficient Ansatz with repetitions equal to 1, 2, and 3, and the solid line is Complex MLP(3,3) with Sign Ansatz, where the total number of parameters in the neural network is 317. Both of them start from the initial state $\ket{+...+}$. The dotted line $99.99\%$ Energy is 0.0001 times the ground state energy.
Table 3: Convergence Efficiency Comparison for the 6 qubit J1-J2 Model
Method Average Optimization Steps Success Probability (%) Summary
sVQNHE 132.67 60 faster convergence with higher reliability
VQE (2-layer HEA) 1676.00 5 more steps and has a lower success rate
Explanation: This table compares the convergence efficiency of sVQNHE and VQE (with a 2-layer hardware-efficient ansatz) for the 6-qubit J1-J2 model. The metrics include the average number of optimization steps required to reach 99.45% of the ground state energy and the probability of success. sVQNHE outperforms VQE by requiring fewer steps (132.67 vs. 1676) and achieving a higher success rate (60% vs. 5%), demonstrating superior efficiency in quantum resource utilization.
Sampling Efficiency Verification
This section presents a compelling argument for performing sampling from the quantum circuit component rather than the classical neural network component that describes the amplitude. Our objective is to verify the computational advantages in the main text argument. As illustrated in Figure 7, we compare the convergence curve of our model with that shown in Figures 3 and 4 of the main text. The key difference is that the set of bitstrings $s$ here is obtained by sampling from the classical neural network component. Compared to the results in Figures 3 and 4 of the main text, we observe that performing sampling from the quantum circuit component is better; fewer shots can yield more stable iterations and faster convergence. This suggests that our quantum circuit can effectively sample from complex probability distributions that are intractable for classical computers, thereby providing a high-quality solution space that closer approximates the true solution space.
Figure 7: Sampling Efficiency Verification under 2D Heisenberg Hamiltonian. The figure illustrates the results obtained under varying numbers of qubits and shots. Sampling is conducted from the classical neural network component. The x-axis denotes the number of iterations, while the y-axis represents the energy. The solid dotted line indicates the ground state energy value of the corresponding model.
Preliminary Resource Consumption Analysis for Larger Instances
Table 4: Resource Comparison for Larger Instances
Qubits Encoding Vertices
brickwork-VQE
measurement cost
Comparison
sVQNHE
measurement cost
17 2 408 3000 $>$ 409
3 2040 24150 $>$ 4081
30 2 1305 9612 $>$ 1306
3 12180 141246 $>$ 24361
Explanation: This table compares the measurement costs of sVQNHE and brickwork-VQE under 2nd and 3rd order polynomial encodings for 17 and 30 qubits. The vertex count represents the maximum number of graph vertices for the maxcut problem corresponding to each qubit count and encoding scheme. Measurement cost reflects the number of measurement circuits required per iteration. sVQNHE consistently requires fewer circuits, demonstrating superior efficiency in quantum resource utilization.
In this section, we compare the measurement requirements for a single iteration of the brickwork-VQE and sVQNHE schemes under 2nd and 3rd order polynomial encodings[36], with equal shots per operator. As shown in the Table. 4, the savings in measurements, defined as the reduction when using sVQNHE compared to brickwork-VQE, scale linearly with the number of vertices in the graph. While this scaling is linear, the absolute savings become significantly large for graphs with many vertices, resulting in a substantial efficiency advantage for sVQNHE in large-scale quantum systems.

### v2

@@ 4. Toward Near-Term Quantum Utility in Large-Scale Combinatorial Optimization
Figure 5: Scaling test of sVQNHE using the MaxCut problem as an example. (a) The relative solution quality for the MaxCut problem, comparing parameter sets 1 and 2 across different problem sizes. The methods sign-VQNHE, brickwork-VQE, and sign-VQE correspond to the sVQNHE method with a 2-layer sign ansatz, the VQE method with brickwork ansatz, and the VQE method with 2-layer sign ansatz, respectively. The depth of brickwork ansatz is determined according to the reference [sciorilli2025towards], that is, the circuit depth is sublinearly related to the number of vertices in the graph. For sign-VQE, the ansatz is the same as the circuit finally generated by the corresponding sVQNHE scheme finally generated, utilizing a two-layer sign ansatz with two diagonal and two non-diagonal layers. The value range of relative solution quality is $[0,1]$. The closer to 1, the better the performance is compared with other algorithms (see Equation (LABEL:eq:R_e) in the Method section for details). For example, in the case of 45 vertices, the quality of the solution of sVQNHE is relatively optimal regardless of whether the parameter set is 1 or 2. (b) The single-step measurement consumption for the same problem sizes. Parameter set 2 features an increased number of parameters in the NN component and doubles the number of iterations compared to parameter set 1. The circuit depth in the brickwork architecture scales sub-linearly with the number of vertices. (c) Exact state-vector simulation for MaxCut on Erdős–Rényi graphs with $m=1485$ vertices. Gray error bars represent one standard deviation. (d) Exact state-vector simulation for MaxClique on Erdős–Rényi graphs with $m=135$ vertices. Gray error bars represent one standard deviation.
@@ Discussion
The results presented in this work support a broader computational principle: hybrid quantum–classical wavefunction learning becomes significantly more scalable when amplitude and phase are trained by different resources rather than by a single quantum circuit. Recent efforts in variational quantum computing have focused on improving expressivity, ansatz structure, or optimizer robustness while still relying on the quantum circuit to learn both components of the wavefunction. Our findings suggest this paradigm may be unnecessarily burdensome for near-term hardware. Demonstrated across a broad range of applications in this study, when the amplitude landscape is learned classically, and the quantum circuit is dedicated solely to learning the phase (sign) structure, optimization becomes measurably easier, gradients remain accessible at larger system sizes, and noise sensitivity decreases without sacrificing expressivity, as evidenced by sVQNHE’s benign variational landscape and gradual transfer mechanism.
Importantly, the performance improvements observed across quantum simulation and combinatorial optimization arise from the division of labor itself. Amplitude learning thrives under classical updates, whereas phase learning benefits from quantum interference in shallow circuits. This explains why sVQNHE’s scaling laws differ from VQE or QAOA: gradient variance no longer collapses exponentially, and measurement cost does not inflate with circuit depth or Hamiltonian complexity. In practice, this enables deeper phase-layer stacking on noisy hardware, with inherent error mitigation under realistic gate noise and finite sampling.
Empirically, in frustrated many-body problems like the 6-qubit J1-J2 model, sVQNHE reduces mean absolute energy error by $98.9\%$ and variance by $99.6\%$ relative to neural baselines, converging $\sim 19\times$ faster than hardware-efficient VQE. Similar gains hold for the 9-qubit 2D Heisenberg model and even sign-free systems like the 9-qubit TFIM or 12-qubit Ising, with error reductions up to $81.1\%$. Extending to molecular systems, sVQNHE yields superior ground-state energies and fidelities for a H2O molecule across geometries. For combinatorial optimization, it excels in MaxCut on Erdős–Rényi graphs, improving solution quality by up to $19\%$ and reducing measurements by $\sim 85\%$ in sampling-based tests up to 45 vertices.
More broadly, these findings shift how hybrid algorithms are conceptualized: the neural network as the primary learning module, while quantum hardware supplying essential but targeted phase transformations. This implies scaling quantum advantage via architectures where quantum roles are small yet critical, as seen in sVQNHE matching the classical Free-Energy Machine on a 1,485-vertex MaxCut instance and a 135-vertex MaxClique instance. Overall, explicit quantum phase learning paired with classical amplitude modeling constitutes a principled hybrid strategy. By refining the quantum hardware’s role, this work reinforces that quantum utility may depend more on complementary classical-quantum designs than on emulating universal circuits.
@@ Methods
(Expressiveness of the Quantum Circuit Ansatz)[larocca2023theory, allcock2024dynamical, holmes2022connecting, kazi2024analyzing, kiani2020learning] Let $\mathfrak{g_{1}}$ and $\mathfrak{g_{2}}$ be the Lie algebras corresponding to the generators $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively, where:
@@
Theorem 2
(Benign Landscape of sVQNHE) For an $l$-layer ($l\geq 1$) sVQNHE algorithm with a commuting diagonal ansatz $W_{l}(\vec{\theta})$, the cost function $\mathcal{L}(\vec{\theta})=\langle\psi(\vec{\theta})|H|\psi(\vec{\theta})\rangle$, where $\psi(\vec{\theta})=Z_{f}F_{l}W_{l}(\vec{\theta})|\phi\rangle=W_{l}(\vec{\theta})|\phi_{f}\rangle$, exhibits at worst polynomially vanishing gradient variance $\text{Var}[\partial_{\theta_{k}}\mathcal{L}]\in\Omega(1/\text{poly}(n))$ if the support set size of $|\phi_{f}\rangle$ is polynomially bounded in $n$. Here, $Z_{f}$ is the normalization coefficient and $|\phi_{f}\rangle=Z_{f}F|\phi\rangle$.
Proof: To prove the theorem, we establish that the variance of the gradient $\partial_{\theta_{k}}\mathcal{L}$ with respect to a parameter $\theta_{k}$ in the $l$-th layer ansatz $W_{l}(\vec{\theta})$ scales at worst polynomially with the number of qubits $n$, avoiding exponential vanishing[cerezo2021cost], due to polynomial constrained input and the commuting diagonal structure.
Since $W_{l}(\vec{\theta})$ is a diagonal unitary composed of commuting Z-related gates, it can be expressed as
$W_{l}(\vec{\theta})=\sum_{x\in\{0,1\}^{n}}e^{i\xi(x,\vec{\theta})}|x\rangle\langle x|,$
where $\xi(x,\vec{\theta})=\sum_{k=1}^{K}\theta_{k}g_{k}(x)$ for basis functions $g_{k}(x)$ (e.g. for the parameters $\theta_{k}$ of gate $R_{zz}$ applied on the $k_{1},k_{2}$ qubit, $g_{k}(y)=(-1)^{y_{k_{1}}+y_{k_{2}}}$, where $y=|y_{1}y_{2}\ldots y_{n}\rangle$).
The gradient is:
$\partial_{\theta_{k}}\mathcal{L}=2\operatorname{Re}\left[\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger}(\vec{\theta}))HW_{l}(\vec{\theta})|\phi_{f}\rangle\right],$
where
$\partial_{\theta_{k}}W_{l}(\vec{\theta})=i\sum_{x}g_{k}(x)e^{i\xi(x,\vec{\theta})}|x\rangle\langle x|,$
so
$\partial_{\theta_{k}}W_{l}^{\dagger}(\vec{\theta})=-i\sum_{x}g_{k}(x)e^{-i\xi(x,\vec{\theta})}|x\rangle\langle x|.$
To assess barren plateaus, compute the variance over random initializations $\vec{\theta}\sim[0,2\pi)^{K}$:
$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]=\mathbb{E}_{\vec{\theta}}[(\partial_{\theta_{k}}\mathcal{L})^{2}]-(\mathbb{E}_{\vec{\theta}}[\partial_{\theta_{k}}\mathcal{L}])^{2}.$
The mean vanishes: $\mathbb{E}_{\vec{\theta}}[\partial_{\theta_{k}}\mathcal{L}]=0$, as $\mathbb{E}_{\vec{\theta}}[e^{i\xi(x,\vec{\theta})}]=0$ for non-constant $\xi$ due to phase cancellation over $0,2\pi)$. Thus, the variance is:
$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]=\mathbb{E}_{\vec{\theta}}[(\partial_{\theta_{k}}\mathcal{L})^{2}].$
Substitute the gradient
$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]=\mathbb{E}_{\vec{\theta}}[\left[\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger})HW_{l}|\phi_{f}\rangle\right]^{2}+\left[\langle\phi_{f}|W_{l}^{\dagger}H(\partial_{\theta_{k}}W_{l})|\phi_{f}\rangle\right]^{2}+2\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger})HW_{l}|\phi_{f}\rangle\langle\phi_{f}|W_{l}^{\dagger}H(\partial_{\theta_{k}}W_{l})|\phi_{f}\rangle].$
Now, substitute the expressions:
$W_{l}(\vec{\theta})=\sum_{x}e^{i\xi(x,\vec{\theta})}|x\rangle\langle x|,\quad\partial_{\theta_{k}}W_{l}(\vec{\theta})=i\sum_{y}g_{k}(y)e^{i\xi(y,\vec{\theta})}|y\rangle\langle y|,\quad\partial_{\theta_{k}}W_{l}^{\dagger}(\vec{\theta})=-i\sum_{y}g_{k}(y)e^{-i\xi(y,\vec{\theta})}|y\rangle\langle y|.$
Then:
$\langle\phi_{f}|(\partial_{\theta_{k}}W_{l}^{\dagger})HW_{l}|\phi_{f}\rangle=-i\sum_{y,z}g_{k}(y)e^{-i\xi(y,\vec{\theta})}\langle\phi_{f}|y\rangle\langle y|H|z\rangle e^{i\xi(z,\vec{\theta})}\langle z|\phi_{f}\rangle,$
$\langle\phi_{f}|W_{l}^{\dagger}H(\partial_{\theta_{k}}W_{l})|\phi_{f}\rangle=i\sum_{y,z}g_{k}(z)e^{-i\xi(y,\vec{\theta})}\langle\phi_{f}|y\rangle\langle y|H|z\rangle e^{i\xi(z,\vec{\theta})}\langle z|\phi_{f}\rangle.$
and
$\langle\phi_{f}|(\partial_{\theta_{k}}V^{\dagger})HV|\phi_{f}\rangle\cdot\langle\phi_{f}|V^{\dagger}H(\partial_{\theta_{k}}V)|\phi_{f}\rangle$
$=\sum_{y_{1},z_{1},y_{2},z_{2}}g_{k}(y_{1})g_{k}(z_{2})e^{-i(\xi(y_{1})-\xi(z_{1}))}e^{i(\xi(y_{2})-\xi(z_{2}))}\langle\phi_{f}|y_{1}\rangle\langle y_{1}|H|z_{1}\rangle\langle\phi_{f}|y_{2}\rangle\langle y_{2}|H|z_{2}\rangle\langle z_{1}|\phi_{f}\rangle\langle z_{2}|\phi_{f}\rangle.$
Averaging over $\vec{\theta}$:
$\mathbb{E}_{\vec{\theta}}\left[e^{-i(\xi(y_{1})-\xi(z_{1})+\xi(y_{2})-\xi(z_{2}))}\right]=\delta_{y_{1}z_{1}}\delta_{y_{2}z_{2}},$
because the expectation over $\vec{\theta}$ cancels unless the phase differences are zero (e.g., $y_{1}=z_{1}=y_{2}=z_{2}$). The same argument applies to other items in $\text{Var}[\partial_{\theta_{k}}\mathcal{L}]$.
Thus we have
$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]\geq\Omega(\sum_{y}g_{k}^{2}(y)|\langle y|H|y\rangle|^{2}|\langle\phi_{f}|y\rangle|^{4}).$
For the form with restricted support set size, where $|\langle\phi_{f}|y\rangle|>0$ only at $poly(n)$ positions—instead of the type where $2^{n}$, by the Cauchy–Schwarz inequality we have:
$\text{Var}[\partial_{\theta_{k}}\mathcal{L}]\geq\Omega(\frac{1}{poly(n)}).$
[…]
@@
$\nabla_{\theta}f(x;\theta)=\frac{1}{2}[f(x;\theta+\frac{\pi}{2})-f(x;\theta-\frac{\pi}{2})].$ (8)
@@
[…]
@@
[…]
@@
In summary, due to the commutation properties of the Sign Ansatz, our algorithm requires fewer measurements compared to general algorithms that rely on parameterized quantum circuits. For instance, considering the 1D J1-J2 model presented below (11), if $|p|\geq 2$ in the Sign Ansatz, the total number of measurements in a single iteration of our algorithm is equal to $3(n_{1}+n_{2})+1$, where $n_{1}$ is the number of $\langle i,j\rangle$, and $n_{2}$ is the number of $\langle\langle i,j\rangle\rangle$.
@@
We use TensorCircuit[zhang2023tensorcircuit] to efficiently obtain all the quantum circuits and the neural network numerical results with the following Ansatz setting, and the details of the VQNHE algorithm are in the TensorCircuit Demo.